OpenStudy (anonymous):

please help solve square root 512k^2 and show work

OpenStudy (anonymous):

it is 512k^2 how do I find the variable and exponent

OpenStudy (anonymous):

\[ \sqrt{512k^2 }=\sqrt{ 16^2 k^2 * 2 } = 4k \sqrt{ 2 }\]

OpenStudy (anonymous):

thanks